OpenStudy (anonymous):

evaluate the series 6+2+2/3+2/9+...a6

OpenStudy (anonymous):

Keep dividing by 3, till you get to the 6th term.

OpenStudy (anonymous):

It appears to be of the form $$\sum_{k}\frac{c}{3^k}$$

OpenStudy (anonymous):

You are given the first four terms, find the fifth and sixth, then add them up.

OpenStudy (anonymous):

how do you find the fifth and sixth terms?

OpenStudy (anonymous):

Each term is a third of the preceding one. A third of 2/9 is about 2/27. A third of that is 2/81. That gives the first six terms in the sequence; just add them up.

OpenStudy (anonymous):

2/27+2/81= 8/81or .0987

OpenStudy (anonymous):

No, you have to add up all six terms. The four they gave you in the problem and the two we just figured out.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

8.98

OpenStudy (anonymous):

That is only an approximation. The exact answer will be a fraction.

OpenStudy (anonymous):

728/81?

OpenStudy (anonymous):

I just did it in my head, but I think that is about right. Do math every day.